# Cherenkov emission by fast-moving uncharged Schwarzschild black hole

### v1

@@ Abstract
We demonstrate that in the presence of an external magnetic field, an uncharged classical Schwarzschild  black hole  moving superluminally in a dielectric with permittivity $\epsilon>1$ produces Cherenkov emission. This is a new physical effect: classical (non-quantum) emission of electromagnetic waves by a completely charge-neutral “particle”. The governing equations (involving General Relativity, electromagnetism, and the physics of continuous media) have no external electromagnetic source - it is the distortion of the initial electromagnetic fields by the gravity of the black hole that plays the role of a superluminally moving source. The effect relies on nonzero values of both the magnetic field and the gravitational radius, as well as on the usual Cherenkov condition on the velocity, $v/c>1/\sqrt{\epsilon}$. Unlike Cherenkov emission by a point charge, the effective source in this case is spatially distributed, with emission generated along the single Cherenkov emission cone. The emitted spectrum is red-dominated, with power $\propto dk_{z}/|k_{z}|$ for wave numbers $|k_{z}|\leq 1/R_{G}$, where $R_{G}$ is the Schwarzschild radius. We comment on possible observability of this processes during black hole-neutron star mergers.
@@ I Introduction
Media with refractive index large than unity are expected in astrophysical settings. For example, strongly magnetized plasmas, with $\omega_{B}\geq\omega_{p}$ ($\omega_{p}$ is plasma frequency, and $\omega_{B}=eB/(m_{e}c)$ is cyclotron frequency) support subluminal normal modes (Barnard and Arons, 1986; Lyutikov, 1999). This is the plasma regime expected, e.g., around various types of neutron stars. During BH-NS mergers, the BH is then moving through magnetized media that supports subluminal waves.
@@ II.1 Electromagnetic emission without an electromagnetic source
Another important difference from the conventional Cherenkov emission is that for a moving BH gravitational pertubations propagates with the speed of light. Conceptually, one can then imagine that the source propagates with the speed of light, but the resulting electromagnetic perturbations are delayed, creating a distributed source of electromagnetic fluctuations.
The relevant references to Cherenkov emission include Tamm and Frank (1937); Jauch and Watson (1948a, b); Bolotovsky and Stolyarov (1974); Bolotovskii (1957). In particular, Refs. Jauch and Watson (1948a, b) discuss Cherenkov emission in the particle frame. This frame has an advantage that the system is stationary. In the conventional Cherenkov case, this leads to the fact that in the particle frame there is only magnetic field (see also Clemmow, 1974). The fields are stationary, but fluctuating with slowly decreasing amplitude. Radiation in the lab frame is then achieved by Lorentz transformation.
@@
Eq. (2) determines the evolution of electromagnetic fields in a dielectric medium in the presence of a moving black hole. It amounts to a generalization to curved space of the flat-space expressions presented in Refs. Jauch and Watson (1948a) [Eq. (14) with gauge Eq. (10)] and Bolotovsky and Stolyarov (1974) [Eqs. (3.4-3.5)]; see also §II.2. In vacuum $\kappa=0$, while in the absence of gravity $g_{\alpha\beta}\to\eta_{\alpha\beta}=\{-1,1,1,1\}$.
@@ II.2 Waves in moving dielectric in flat space
Equation (6) has the fluctuating part of the field “on the left” and the source part “on the right”. The operator $\hat{L}_{0}$ acting on the fluctuating part is none other than the wave operator in a uniformly moving dielectric, in the absence of any gravitational effects (see §II.4 and Appendix C for discussion of the concept of“uniformly moving” in the case of curved space). Its structure has been discussed in Refs. Jauch and Watson (1948a, b); Bolotovskii (1957). Here, we list a few of its properties that we will need in the following.
@@
$\cos^{2}\theta=\frac{1}{\beta^{2}\gamma^{2}\kappa}$ (20)
(the velocity now refers to that of the medium in the BH’s frame). This is the Cherenkov cone in the BH frame.
@@ II.4 Choice of the velocity field
We also need to impose the velocity structure $u^{\mu}$. In our approach, we are not solving for the dynamics of matter. Qualitatively, the medium is assumed to be a massless ether-like substance that just modifies the properties of electromagnetic waves. Motion of the medium is important, though: in a moving dielectric electromagnetic energy both propagates and is advected by the bulk motion.
@@ II.5 Stationary perturbation in BH frame
There is then a subtlety: in terms of Fourier transforms with respect to $z$, that means that we must now choose between the outgoing or incoming asymptotic of the fields at infinity. The outgoing one corresponds to the Cherenkov radiation. This implies, implicitly, some time-dependence, namely, that the source was turned-on at some time in the far past. (This relates, e.g., to the Landau rule Landau and Lifshitz (1960)).
@@ III.1.1 Source choice I
The magnetic field corresponding to (32) is (Fig. 1)
$B_{\rho}=\frac{M\rho z}{\left(\rho^{2}+z^{2}\right)^{3/2}}\times\left(1+\frac{M}{2\sqrt{\rho^{2}+z^{2}}}\right)^{3}B_{0}\approx\frac{M\rho z}{\left(\rho^{2}+z^{2}\right)^{3/2}}B_{0}\mbox{}$
$B_{z}=\left(1+\frac{M\left(z^{2}-\rho^{2}\right)}{2\left(\rho^{2}+z^{2}\right)^{3/2}}\right)\times\left(1+\frac{M}{2\sqrt{\rho^{2}+z^{2}}}\right)^{3}B_{0}\approx\left(1+\frac{M\left(\rho^{2}+2z^{2}\right)}{\left(\rho^{2}+z^{2}\right)^{3/2}}\right)B_{0}$ (33)
where the final expressions retain only the leading corrections in the $M\to 0$ limit.
@@ III.2 Main calculations: parallel propagation
Figure 2: Vector potential $|\tilde{a}|$, Eq. (39), Choice I. The $\Lambda=0$ curve is given by (40) - it clearly serves as a separatrix between regular and Cherenkov regimes.
@@
$|\tilde{a}|\propto\frac{e^{\pm i|\Lambda|k\rho}}{\sqrt{k\rho}}$ (42)
@@
The term in $A^{(1)}_{\phi}$ responsible for radiation in the Cherenkov regime is then
$A_{\phi}=\frac{4\Lambda}{1-\Lambda^{2}}\times\frac{MB_{0}}{k}\rho K_{1}(\Lambda k\rho)$ (43)
where $\Lambda\to\pm i|\Lambda|$.
It is also instructive to compare the effect of the remaining (non-Cherenkov) terms in (39) for our two different choices of the source. We will do that by going over to the real space in the non-Cherenkov regime $\Lambda^{2}>0$. For Choice II, upon multiplication by $\rho$ and the Fourier transform to the coordinate space, we obtain
@@
Observe that in the absence of the medium, i.e., for $\Lambda\to 0$, when expanded to the first order in $M$, (44) becomes
$A_{\phi}(\rho,z)\approx\frac{1}{2}B_{0}\rho^{2}+\frac{MB_{0}\rho^{2}}{(\rho^{2}+z^{2})^{1/2}}\,.$ (45)
The second (correction) term in this expression is precisely reproduced by the $\Lambda\to 1$ limit of the exact solution (32). Thus, the same term that is included as a part of the zeroth-order approximation for Choice I, appears as a part of the first-order solution for Choice II.
@@ IV.1 The source
In the black hole frame there are now crossed electric and magnetic fields at infinity. We chose the velocity of the black hole along $z$ and the asymptotic magnetic field along $x$, so that the asymptotic electric field is along $-y$. The isotropic Cartesian coordinates all have the same diagonal elements, equal to $g_{zz}$ in Eq. (II.3).
@@
We have already demonstrated that for parallel propagation the two choices result in the same radiative term. As it turns out, computations for the choice (52) are somewhat simpler, so we proceed with it here. For reference, the sources corresponding to (51) are given in Appendix B.
In this section, we choose the 4-velocity of the medium in the form (25). The corresponding expressions for a more general (one-parameter) choice are given in Appendix C.
@@ IV.2 Fluctuating components and Poynting power for perpendicular propagation
For perpendicular propagation the source separates into three parts : $S_{y}$, $S_{0}$ and $S_{z}$. Each component will produce correspondingly three fluctuating potenetials $A_{y}$, $A_{0}$ and $A_{z}$. In the BH frame the corresponding electromagnetic fields are stationary. Boosting to the lab frame, we then obtain propagating waves. Importantly, even though each fluctuating component can be calculated independently, the resulting Poynting flux includes also cross-terms.
@@
${\cal P}_{\perp}=2\beta\times\left(\frac{|\Lambda|^{2}}{1+|\Lambda|^{2}}\right)^{2}\times\left(\cos^{2}(\phi)+\frac{\beta^{2}\sin^{2}(\phi)}{\epsilon}\right)\times\frac{dk}{2\pi k}\times(\gamma MB_{0})^{2}\mbox{}$
@@ V Observability
A simple order-of-magnitude estimate of Cherenkov power is given by Eqns. (50) and (66). Returning to dimensional quantities
@@ VI Discussion
(see Appendix §C for an interesting exception).
Scaling (72) indicates that the effect is the interplay between gravity $(GM)$, magnetic field $B_{0}$, and physics of movement of continuous media $\Lambda$: $\kappa\neq 0$, $\beta\neq 0$ is needed, plus the Cherenkov condition.
The process involves an number of subtle physical effects:
@@ Appendix A Relevant mathematical formulae
In many cases we deal with equation
@@
The following are relevant solutions of the inhomogeneous equation (it appears that the following relations are not listed in the standard textbooks on Bessel functions)
@@
$S=K_{1}(\rho_{1})e^{i\phi}tof=\left(K_{1}(\rho_{1})-\Lambda K_{1}(\Lambda\rho_{1})\right)\frac{e^{i\phi}}{1-\Lambda^{2}}\mbox{}$
@@
Appendix B Sources for perpendicular motion with initial state (51)
Perturbations of both the electric field and magnetic field will each produce Cherenkov-like emission (in some cases adding destructively).
Following the procedure described above, we find the total source
$S^{0}=-\left(\frac{3\beta^{2}\gamma^{2}\rho^{2}}{\left(\rho^{2}+z^{2}\right)^{5/2}}+\frac{2}{\left(\rho^{2}+z^{2}\right)^{3/2}}\right)\beta\rho\sin\phi\times\kappa{\cal B}_{0}M\mbox{}$
$S^{x}=0\mbox{}$
$S^{y}=\frac{z\left(2z^{2}-\rho^{2}\right)}{\left(\rho^{2}+z^{2}\right)^{5/2}}\beta^{2}\gamma^{2}\times\kappa{\cal B}_{0}M\mbox{}$
$S^{z}=\frac{\rho\left(\rho^{2}+4z^{2}\right)}{\left(\rho^{2}+z^{2}\right)^{5/2}}\beta^{2}\gamma^{2}\sin\phi\times\kappa{\cal B}_{0}M$ (78)
where we used $y=\rho\sin\phi$. (Component $S_{x}=0$ is due to interference of E-produced and B-produced terms, each is non-zero independently).
The Fourier transform of the source is
$S^{0}=-2k_{z}^{2}\beta\left(\beta^{2}\gamma^{2}\rho_{1}K_{2}(\rho_{1})+2K_{1}(\rho_{1})\right)\sin\phi\times\cos\left(k_{z}z\right)\times\beta{\cal B}_{0}\kappa M\mbox{}$
$S^{x}=0\mbox{}$
$S^{y}=2\beta^{2}\gamma^{2}k_{z}\left(\rho_{1}K_{1}(\rho_{1})-2K_{0}(\rho_{1})\right)\times\sin\left(k_{z}z\right)\times\beta{\cal B}_{0}\kappa M\mbox{}$
$S^{z}=2\beta^{2}\gamma^{2}k_{z}^{2}\left(4K_{1}(\rho_{1})-\rho_{1}K_{2}(\rho_{1})\right)\sin\phi\times\cos\left(k_{z}z\right)\times\beta{\cal B}_{0}\kappa M$ (79)
Explicit sine and cosine transforms are indicated, and overall normalization ×$\times{\cal G}$ has been omitted. We also explicitly give the source for the $0-z$ combination. (The $S_{y}$ source is weakly divergent at $\rho_{1}\to 0$ as $\ln\rho_{1}$.)
Appendix C Choices of velocity field for perpendicular propagation

### v4

# Cherenkov emission by a fast-moving uncharged Schwarzschild black hole

@@ Abstract
We demonstrate that, in the presence of an external magnetic field, an uncharged classical Schwarzschild  black hole  moving superluminally in a dielectric with permittivity $\epsilon>1$ produces Cherenkov emission. This is a new physical effect: classical (non-quantum) emission of electromagnetic waves by a completely charge-neutral “particle.” The governing equations (involving general relativity, electromagnetism, and the physics of continuous media) have no external electromagnetic source—it is the distortion of the initial electromagnetic fields by the gravity of the black hole that plays the role of a superluminally moving source. The effect relies on nonzero values of both the magnetic field and the gravitational radius, as well as on the usual Cherenkov condition on the velocity, $v/c>1/\sqrt{\epsilon}$. Unlike Cherenkov emission by a point charge, the effective source in this case is spatially distributed, with emission generated along the single Cherenkov emission cone. The emitted spectrum is red-dominated, with power $\propto dk_{z}/|k_{z}|$ for wave numbers $|k_{z}|\leq 1/R_{G}$, where $R_{G}$ is the Schwarzschild radius. We comment on possible observability of this process during black hole–neutron star mergers.
@@ I Introduction
Media with refractive index larger than unity are expected in astrophysical settings. For example, strongly magnetized plasmas, with $\omega_{B}\geq\omega_{p}$ ($\omega_{p}$ is plasma frequency, and $\omega_{B}=eB/(m_{e}c)$ is cyclotron frequency) support subluminal normal modes (Barnard and Arons, 1986; Lyutikov, 1999). This is the plasma regime expected, e.g., around various types of neutron stars. During BH-NS mergers, the BH is then moving through magnetized media that supports subluminal waves.
@@ II.1 Electromagnetic emission without an electromagnetic source
Another important difference from the conventional Cherenkov emission is that for a moving BH gravitational perturbations propagate with the speed of light. Conceptually, one can then imagine that the source propagates with the speed of light, but the resulting electromagnetic perturbations are delayed, creating a distributed source of electromagnetic fluctuations.
The relevant references to Cherenkov emission include Tamm and Frank (1937); Jauch and Watson (1948a, b); Bolotovskii and Stolyarov (1975); Bolotovskii (1957). In particular, Refs. Jauch and Watson (1948a, b) discuss Cherenkov emission in the particle frame. This frame has an advantage that the system is stationary. In the conventional Cherenkov case, this leads to the fact that in the particle frame there are stationary electric and magnetic fields (see also Clemmow (1974)). These fields oscillate in space with slowly decreasing amplitudes. Radiation in the lab frame is then achieved by Lorentz transformation.
@@
Eq. (2) determines the evolution of electromagnetic fields in a dielectric medium in the presence of a moving black hole. It amounts to a generalization to curved space of the flat-space expressions presented in Refs. Jauch and Watson (1948a) [Eq. (14) with gauge Eq. (10)] and Bolotovskii and Stolyarov (1975) [Eqs. (3.4)–(3.5)]; see also §II.2. In vacuum $\kappa=0$, while in the absence of gravity $g_{\alpha\beta}\to\eta_{\alpha\beta}=\{-1,1,1,1\}$.
@@ II.2 Waves in moving dielectric in flat space
Equation (6) has the fluctuating part of the field “on the left” and the source part “on the right”. The operator $\hat{L}_{0}$ acting on the fluctuating part is none other than the wave operator in a uniformly moving dielectric, in the absence of any gravitational effects (see §II.4 and Appendix B for discussion of the concept of“uniformly moving” in the case of curved space). Its structure has been discussed in Refs. Jauch and Watson (1948a, b); Bolotovskii (1957). Here, we list a few of its properties that we will need in the following.
@@
$\cos^{2}\theta=\frac{1}{\beta^{2}\gamma^{2}\kappa}=\frac{1}{1-\Lambda^{2}}=\frac{1}{1+|\Lambda|^{2}}$ (20)
(the velocity now refers to that of the medium in the BH’s frame). This is the Cherenkov cone in the BH frame (for $\Lambda^{2}<0$).
@@ II.4 Choice of the velocity field
We also need to impose the velocity structure $u^{\mu}$. In our approach, we are not solving for the dynamics of matter. Motion of the medium is important, though: in a moving dielectric electromagnetic energy both propagates and is advected by the bulk motion.
@@ II.5 Stationary perturbation in BH frame
There is then a subtlety: in terms of Fourier transforms with respect to $z$, that means that we must now choose between the outgoing or incoming asymptotic of the fields at infinity. The outgoing one corresponds to the Cherenkov radiation. Thus, there is implicitly some time-dependence, namely, that the source was turned-on at some time in the far past. (This relates, e.g., to the Landau rule Landau and Lifshitz (1960)).
@@ III.1.1 Source choice I
The components of the magnetic field corresponding to (32) and computed according to (Landau and Lifshitz, 1975, §90) are (Fig. 1)
$B^{\rho}=\frac{M\rho z}{\left(\rho^{2}+z^{2}\right)^{3/2}}\times\left(1+\frac{M}{2\sqrt{\rho^{2}+z^{2}}}\right)^{-3}B_{0}\approx\frac{M\rho z}{\left(\rho^{2}+z^{2}\right)^{3/2}}B_{0}\mbox{}$
$B^{z}=\left(1+\frac{M\left(z^{2}-\rho^{2}\right)}{2\left(\rho^{2}+z^{2}\right)^{3/2}}\right)\times\left(1+\frac{M}{2\sqrt{\rho^{2}+z^{2}}}\right)^{-3}B_{0}\approx\left(1-\frac{M\left(2\rho^{2}+z^{2}\right)}{\left(\rho^{2}+z^{2}\right)^{3/2}}\right)B_{0}$ (33)
@@ III.2 Main calculations: parallel propagation
Figure 2: Vector potential $|\tilde{a}|$ (excluding normalization $MB_{0}/k$), Eq. (39), Choice I. The $\Lambda=0$ curve is given by (40) - it clearly serves as a separatrix between regular and Cherenkov regimes.
@@
$\tilde{a}\propto\frac{e^{\pm i|\Lambda|k\rho}}{\sqrt{k\rho}}$ (42)
@@
It is also instructive to go over to the coordinate space. We will do that starting with Eq. (39) for Choice II. In the non-Cherenkov regime, $\Lambda^{2}>0$, the first-order deformation $A^{(1)}_{\phi}(\rho,z)$ is obtained directly by multiplying (39) by $\rho$ and Fourier transforming. The result is
@@
Observe that in the absence of the medium, i.e., for $\Lambda\to 1$, Eq. (43) becomes
$A^{(1)}_{\phi}(\rho,z)=\frac{MB_{0}\rho^{2}}{(\rho^{2}+z^{2})^{1/2}}\,.$ (44)
This precisely reproduces the $O(M)$ term in the expansion of the exact solution (32). Thus, the same term that is included as a part of the zeroth-order approximation for Choice I, appears as a part of the first-order solution for Choice II.
Observe also that, close to the boundary with the Cherenkov regime, $\Lambda^{2}\to 0$, the second term in the bracket in (43) approaches $|z|$ (at a given finite $\rho$), meaning that the magnetic field undergoes a rapid charge near the $z=0$ plane—a precursor to the formation of the Cherenkov leading shock-like wavefront.
In the Cherenkov regime, $\Lambda^{2}<0$, the term in $A^{(1)}_{\phi}(\rho,k_{z})$ responsible for Cherenkov radiation is
$A_{\phi}^{(rad)}(\rho,k_{z})=\frac{4\Lambda}{1-\Lambda^{2}}\times\frac{MB_{0}}{k}\rho K_{1}(\Lambda k\rho)\,.$ (45)
Going over to the coordinate space requires analytically continuing this to $\Lambda=\pm i|\Lambda|$ (using the procedure described in Subsec. III.3 below) and Fourier transforming the result. We find
$A_{\phi}^{(rad)}(\rho,z)=-\frac{2MB_{0}}{1+|\Lambda|^{2}}\left\{\begin{array}[]{lr}z,&|z|<|\Lambda|\rho\\ z+(1-\mbox{sgn\,}z)\sqrt{z^{2}-|\Lambda|^{2}\rho^{2}},&|z|\geq|\Lambda|\rho\end{array}\right.$ (46)
where sgn $\mbox{sgn\,}z$ denotes the sign of $z$. This potential is smooth for $z>0$ (i.e., in front of the black hole) but exhibits a shock-like wavefront at $z=-|\Lambda|\rho$. The direction normal to this wavefront coincides with that of Cherenkov emission, Eq. (20).
@@ IV.1 The source
In the black hole frame there are now crossed electric and magnetic fields at infinity. We chose the velocity of the black hole along $z$ and the asymptotic magnetic field along $x$, so that the asymptotic electric field is along $y$ (Fig. 3). The isotropic Cartesian coordinates all have the same diagonal elements, equal to $g_{zz}$ in Eq. (II.3).
@@
We have already demonstrated that for parallel propagation the two choices result in the same radiative term. As it turns out, computations for the choice (53) are somewhat simpler, so we proceed with it here.
In this section, we choose the 4-velocity of the medium in the form (25). The corresponding expressions for a more general (one-parameter) choice are given in Appendix B.
@@ IV.2 Fluctuating components and Poynting power for perpendicular propagation
For perpendicular propagation the source separates into three parts : $S_{y}$, $S_{0}$ and $S_{z}$. Each component will produce correspondingly three fluctuating potentials $A_{y}$, $A_{0}$ and $A_{z}$. In the BH frame the corresponding electromagnetic fields are stationary. Boosting to the lab frame, we then obtain propagating waves. Importantly, even though each fluctuating component can be calculated independently, the resulting Poynting flux includes also cross-terms.
@@
${\cal P}_{\perp}=2\beta\times\left(\frac{|\Lambda|^{2}}{1+|\Lambda|^{2}}\right)^{2}\times\left(\cos^{2}(\phi)+\frac{\beta^{2}\sin^{2}(\phi)}{\epsilon}\right)\times\frac{dk_{z}}{2\pi k}\times(\gamma MB_{0})^{2}\mbox{}$
@@ V Observability
A simple order-of-magnitude estimate of Cherenkov power $L$ is given by Eqns. (51) and (67). Returning to dimensional quantities
@@ VI Discussion
(see Appendix §B for an interesting exception).
Scaling (73) indicates that the effect is the interplay between gravity $(GM)$, magnetic field $B_{0}$, and physics of movement of continuous media: both $\kappa\neq 0$, $\beta\neq 0$ are needed, plus the Cherenkov condition.
The process involves a number of subtle physical effects:
@@ Appendix A Relevant mathematical formulae
In many cases we deal with the equation
@@
The following are the relevant solutions of the inhomogeneous equation (note the ones for $S=\rho_{1}Z(\rho_{1})e^{im\phi}$, which do not seem to be listed in the standard textbooks on Bessel functions)
@@
$S=K_{1}(\rho_{1})e^{i\phi}\to f=\left(K_{1}(\rho_{1})-\Lambda K_{1}(\Lambda\rho_{1})\right)\frac{e^{i\phi}}{1-\Lambda^{2}}\mbox{}$
@@
Appendix B Choices of velocity field for perpendicular propagation